1.) Consider the hypothetical market for taking the TB test shown below for Parry County, Alabama. Answer based on the graph provided.MSB $30 $20 S 1 200 2000 20.000 -$20 D = MPB

A.)How many people with get the test if it is not subsidized (marginal cost is $20)?

B.)Is this free-market outcome optimal? Why (not)?

C.)How many people will get the test if the government makes it free? What would that cost the government?

D.)How many people will get the test if they get paid $20? How much does the $20 incentive program cost the government (subsidy + cost of test)?

E.)Can you relate this to COVID-19?
Answer #1

A) at eqm, P = MC = 20

So Q*= 200

Where MC = PMB

so 200 people will get test positive

B) free market outcome is not pareto Optimal , bcoz at P = 20, MSB curve cuts MC at relatively higher number of individuals

Thus market outcome is less than socially optimal level

C) if govt makes it free, then PMB = demand D cuts X axis at 2000

2000 people will get the test

cost = 2000*20= $ 40,000

D) with Subsidy, P = -20, then Q= 20,000

Cost to govt = (20+20)*20,000

= 80,000

E) again in Corona virus case, the same issue arrives, when there is a Fixed price of test, people are dis-incenticized to take the test

If it's made free or subsidy is provided, then number of people, who get the test rises